OpenStudy (anonymous):

Is the given answer a reasonable solution to the problem? 8.89 • 12.13 = 107.8357 A. Yes B. No

OpenStudy (anonymous):

and I also need help on this one.

OpenStudy (anonymous):

Is this practice in estimating? I'm not sure what the goal of this question is

OpenStudy (anonymous):

for which one?

OpenStudy (anonymous):

The question of the thread.

OpenStudy (anonymous):

I'm not really sure. I haven o idea on how to do this.:p

OpenStudy (anonymous):

first is A

OpenStudy (anonymous):

Well, the most straight forward way is to just do the multiplication. \[12.13\times 8.89 = ?\]

OpenStudy (anonymous):

second is C d=rt finding t? then divide d by r = t

OpenStudy (anonymous):

When did this place become "Here's the answer?"

OpenStudy (anonymous):

d=rt t=d/r do the same thing as above, just opposite. m=k/s k=sm <---

OpenStudy (anonymous):

or k=ms

OpenStudy (minecrafter64):

I think its D